# Anant-Net: Breaking the Curse of Dimensionality with Scalable and Interpretable Neural Surrogate for High-Dimensional PDEs

### v2

@@ 3.2 High-dimensional sine-Gordon equation
Dimension Architecture Optimizer ##\##Collocation ##\##Data
@@ 3.3 High-dimensional Allen-Cahn equation
Dimension Architecture Optimizer ##\##Collocation ##\##Data
@@
3.4 Effect of Boundary Data Sampling on Anant-Net Accuracy
@@
[…]
@@ 3.5 Anant-KAN Results
Figure 13: Point-wise absolute testing error for random 3D hyper-surfaces of 21D, 99D (left) Poisson equation (center) sine-Gordon equation, (right) Allen-Cahn equation. Here, ”$\mathsf{ActiveDims}$: $(p,q,r)$” represents the active dimensions in the $d$-dimensional space where gradients are non-zero and $\mathrm{D}-\{p,q,r\}$ are inactive. Also, $\mathrm{D}=\{1,2,\dots,d\}$ is a set which constitutes all the dimensions.
@@ 4 Conclusions
In this work, we introduced Anant-Net, a novel and scalable neural architecture tailored for solving high-dimensional partial differential equations (PDEs) defined over hypercubic domains. Traditional numerical and learning-based methods often struggle with the curse of dimensionality, especially when the side length of the domain $L>1$, resulting in significant computational bottlenecks. Anant-Net overcomes these limitations through several key innovations, summarized below:
@@
  • The model demonstrates the ability to solve both linear and nonlinear PDEs in up to 300 dimensions within a few hours on a single GPU. We report the mean and standard deviation (variance) of relative $L_{2}$ testing error on randomly selected test points from the high-dimensional spaces. Furthermore, the accuracy can be enhanced by increasing the number of training and collocation points.
@@ Appendix A Universal Approximation Theorem: Through the lens of Anant-Net
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$, we can find a sufficiently large $r>0$ and neural networks $f_{j}$ and $g_{j}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (12)
@@
where, ^={^$\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, ^$\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X\text{\textbackslash}\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y\text{\textbackslash}\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}$.
Proof. Let ${\phi_{i}}$ and ${\psi_{j}}$ be the orthonormal basis for $L^{2}(X)$ and $L^{2}(Y)$ respectively. Then ${\phi_{i}\psi_{j}}$ forms the orthonormal basis for $L^{2}(X\times Y)$. Therefore, we can find a sufficiently large r, consider an active dimension set $\mathcal{B}_{k}$ that results in the solution $u_{k}$ such that,
@@
[…]
@@
$\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\sum_{k}^{|\hat{\mathcal{B}}|}\varepsilon=|\hat{\mathcal{B}}|\cdot\varepsilon$
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}b_{i}\tilde{f_{i}}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}\{b_{i}\tilde{f_{i}}\}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}=\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}t_{i}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{r}f_{i}g_{i}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@ Appendix B Hyperparameters for Anant-Net
The main hyperparameters for Anant-Net include: the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), number of boundary/data grids (#$\#BG$), and the sampling frequency ($S_{\text{FREQ}}$). The hyperparameters used for the main results presented in this paper are discussed in this section. For clarity, the tables presented below follow the format: ##\##Collocation $=(N_{C}\times N_{C}\times N_{C})\times\#CG$ and ##\##Data $=(N_{B}\times N_{B}\times N_{B})\times\#BG$ for the current cases dealing with 3 body networks. However, if any future work deals with 4 body networks, then ##\##Collocation $=(N_{C}\times N_{C}\times N_{C}\times N_{C})\times\#CG$ and ##\##Data $=(N_{B}\times N_{B}\times N_{B}\times N_{B})\times\#BG$ and likewise.
@@
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
@@
Dimension ##\##Collocation ##\##Data
@@
Table 9: Details on collocation and data points sampled for cases in Table 8 for solving high-dimensional Poisson equation
@@
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
@@
Dimension ##\##Collocation ##\##Data
@@
Table 11: Details on collocation and data points sampled for cases in Table 10 for solving high-dimensional sine-Gordon equation
@@
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
@@
Dimension ##\##Collocation ##\##Data
@@
Table 13: Details on collocation and data points sampled for cases in Table 12 for solving high-dimensional Allen-Cahn equation
@@
Equation Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
@@
Table 14: Anant-KAN architecture used for solving high-dimensional Poisson equation, sine-Gordon equation and Allen-Cahn equation. Note that Spline basis is used for all the cases with polynomial order ($k$) = 2 and grid size ($G$) = 5
Equation Dimension ##\##Collocation ##\##Data

### v3

@@ 3.2 High-dimensional sine-Gordon Equation
Dimension Architecture Optimizer #\#Collocation #\#Data
@@ 3.3 High-dimensional Allen-Cahn Equation
Dimension Architecture Optimizer #\#Collocation #\#Data
@@
3.4 Effect of Boundary Data Sampling, Collocation Points Sampling, and Batch Size on Anant-Net Performance
@@
Figure 9: Testing accuracy on random samples and training run-time for 300D (Left) Poisson equation and (Right) sine-Gordon equation.
@@
Figure 10: Sensitivity of (Left) collocation grid density ($CG\times(N_{C})^{B}$) and (Right) batch size ($B$) on testing accuracy and training run-time for 300D sine-Gordon equation while keeping all the other hyperparameters (such as the sampling frequency, depth/width of the body networks etc.) fixed.
The sensitivity of hyperparameters are further studied from the perspective of collocation grids and batch size ($i.e.$ number of body networks). This is shown in the Figure 10. As the number of collocation samples increases, the accuracy of Anant-Net increases with negligible changes in the training time as shown in Figure 10 (Left). Note that number of collocation samples is increased by increasing the resolution of the collocation grids ($N_{C}$) while keeping the number of collocation grids (#CG\#CG) constant. The increase in the resolution of collocation grids is seamlessly handled by the tensor product operation without any memory bottleneck while keeping the GPU below its saturation limit leading to constant runtime as shown in the Figure 10 (Left). Figure 10 (Right) shows the sensitivity of batch size ($B$) on the performance of Anant-Net while keeping all the other hyperparameters unchanged and maintaining almost similar number of collocation and data samples across all the cases. Here, we recognize batch size = 3 as the optimal batch size for the 300D sine-Gordon equation and reducing the batch size to 2 showcases dramatic drop in the accuracy of Anant-Net on randomly selected testing data samples. This shows that batch size = 2 does not have sufficient capacity to approximate the solution to a 300D non-linear partial differential equation. However, when batch size = 4, although the accuracy is substantially better than batch size = 2, we observe sufficient drop in the accuracy of Anant-Net compared to batch size = 3. Although this seems like an anomaly at first, it is worthwhile to note that the accuracy of Anant-Net for batch size = 4 can be improved further by fine-tuning the fixed hyperparameters (such as the sampling frequency, depth/width of the body networks etc.). The difference in run-time observed across different values of batch size in Figure 10 (Right) is representative of the difference in number of collocation grids (#CG\#CG) and number of boundary data grids (#BG\#BG) which are needfully changed in order to keep the total number of collocation samples ($CG\times(N_{C})^{B}$) and total number of data samples ($BG\times(N_{B})^{B}$) constant across the different cases of varying batch size ($B$).
@@ 3.5 Anant-KAN Results
Figure 14: Point-wise absolute testing error for random 3D hyper-surfaces of 21D, 99D (left) Poisson equation (center) sine-Gordon equation, (right) Allen-Cahn equation. Here, $\mathsf{ActiveDims}$: (p,q,r)(p,q,r)” represents the active dimensions in the $d$-dimensional space where gradients are non-zero and $\mathrm{D}-\{p,q,r\}$ are inactive. Also, $\mathrm{D}=\{1,2,\dots,d\}$ is a set which constitutes all the dimensions.
3.6 High-dimensional Transient Heat Equation
The high-dimensional heat equation is a parabolic PDE that arises in various scientific and engineering disciplines, which governs unsteady processes such as thermal diffusion, mass diffusion and the like. Here, we consider the heat equation of the following form:
$\partial_{t}u-\Delta u=f(x,t),(x,t)\in\Omega\times[0,T],$ (12)
$u(x,t)=g(x,t),(x,t)\in\partial\Omega\times[0,T],$ (13)
$u(x,0)=h(x),x\in\Omega,$ (14)
where $\Omega\in[-1,1]^{d}$, $T=1$ and the exact solution is assumed to be of the form:
$u_{\text{exact}}(x)=\cos\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}\exp(-t),$
and
$f(x)=\bigl{(}\frac{1}{d}-1\bigr{)}\cos\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}\exp(-t),$
$g(x,t)=\cos\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}\exp(-t),$
$h(x)=\cos\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}.$
Dimension Architecture Optimizer #\#Collocation #\#Data
20D $\bigl{[}1,64,64,15\bigr{]}\times 1,\bigl{[}10,64,64,15\bigr{]}\times 2$ Adam 54880 50000
100D $\bigl{[}1,64,64,50\bigr{]}\times 1,\bigl{[}50,64,64,50\bigr{]}\times 2$ Adam 164640 90000
300D $\bigl{[}1,64,64,160\bigr{]}\times 1,\bigl{[}150,64,64,160\bigr{]}\times 2$ Adam 411600 164000
Table 6: Anant-Net architecture used for solving high-dimensional heat equation. Note that #\#Data represents the number of data points. In this case, #\#Data == Number of boundary samples == Number of initial samples.
Figure 15: Heat Equation: (Left) Training time profiling; (Right) mean testing error, for the 20D, 100D, and 300D cases with the additional time ($t$) dimension on T4 GPU.
Figure 16: Heat Equation: Training loss for the 20D, 100D, and 300D cases with the additional time ($t$) dimension on T4 GPU.
In this case we employ layer-wise adaptive nonlinear activation function (sin) with trainable adaptive parameters for all body networks. We observe consistent scaling for run time and exceptional accuracy for Anant-Net while approximating the solution for the parabolic, high-dimensional, non-homogeneous heat equation for 20D, 100D and 300D. The performance of Anant-Net is shown in Figure 15 and Table 6 provides more information about the architecture, number of collocation samples and data samples used for training. It is worthwhile to note that, unlike in the previous cases that dealt with a steady state high-dimensional PDE, here we deal with an unsteady high-dimensional PDE leading to the presence of the time derivative along with other spatial derivatives in the formulation of the partial differential equation. To this end, we assign an entire body network for learning the dynamics of the high-dimensional system or in other words, the time dimension ($t$) is always an active dimension for that body network. However, the other spatial dimensions ($x_{i}$) are distributed to the remaining body networks and these body networks will have simultaneous active and inactive dimensions at every iteration during the training process which are randomly sampled like in the previous sections. The convergence behavior for the heat equation for 20D, 100D and 300D are shown in Figure 16. More information about the architecture and hyperparameters are provided in APPENDIX B.
Figure 17: Heat Equation: (Left) Training time profiling; (Right) mean testing error, for the 100D with the additional time ($t$) dimension for Anant-Net (on T4 GPU) and Anant-KAN (on High RAM CPU) architectures. For Anant-KAN, we use the spline basis with order of basis $(k)=2$ and grid size $(G)=5$.
Figure 18: Point-wise absolute testing error for random 3D hyper-surfaces of 100D (Top) Anant-Net (Bottom) Anant-KAN, for Heat equation. Here, $\mathsf{ActiveDims}$: (p,q,r)(p,q,r)” represents the active dimensions in the $d$-dimensional space where gradients are non-zero and $\mathrm{D}-\{p,q,r\}$ are inactive. Also, $\mathrm{D}=\{1,2,\dots,d\}$ is a set which constitutes all the dimensions. Since, the problems involves time derivative, a body network is dedicated towards learning the dynamics. Hence, $p=\text{time }(t)$ is fixed for all of the above cases.
We also perform an empirical study for the current case using Anant-KAN in order to show its capability for solving high-dimensional parabolic PDEs with time derivatives. As it was the case for Anant-Net, we dedicate a body network of Anant-KAN exclusively for learning the dynamics of the high-dimensional PDE and time is always an active dimension for this body network. However, all the other dimensions in the high-dimensional space are distributed to rest of the body networks in Anant-KAN and the dimensions are simultaneously active or inactive for these body networks. Similar to Anant-Net, Anant-KAN also showcases good accuracy for 100D non-homogeneous heat equation as shown in Figure 17. Also note that Anant-KAN was run on a High-RAM CPU as opposed to the T4 GPU since the implementation of Kolmogorov Arnold Networks in the high-dimensional setting poses a major memory bottleneck. Although, currently we overcome this by switching to a high RAM CPU configuration, the implementation can be optimized further to overcome the memory bottleneck and this may be considered for a future work.
@@ 4 Conclusions
High-dimensional partial differential equations (PDEs) arise in a broad range of scientific and engineering applications, including uncertainty quantification, plasma physics, quantitative finance, molecular dynamics, gas-kinetic theory, etc. The ability to efficiently solve such PDEs is important for advancing simulation, analysis, and informed decision-making in complex, real-world systems. In this work, we introduced Anant-Net, an efficient and scalable neural architecture tailored for solving high-dimensional PDEs defined over hypercubic domains. Traditional numerical and learning-based methods often struggle with the curse of dimensionality, especially when the side length of the domain $L>1$, resulting in significant computational bottlenecks. Anant-Net overcomes these limitations through several key innovations, summarized below:
@@
  • The model demonstrates the ability to solve both linear and nonlinear, steady state and transient PDEs up to 300 dimensions within a few hours on a single GPU. We report the mean and standard deviation (variance) of relative $L_{2}$ testing error on randomly selected test points from the exponentially increasing high-dimensional spaces. Furthermore, the accuracy can be enhanced by increasing the number of training and collocation points.
@@
Some of the known limitations of Anant-Net are as follows: Anant-Net is primarily designed for high-dimensional PDEs in Euclidean domains where a tensor product structure is present, enabling efficient computations with complexity scaling as $\mathcal{O}(Nd)$. This structure allows straightforward handling of linear additive operators such as the Laplacian, where contributions from individual dimensions are independent and additive. However, for more complex operators, like variable-coefficient diffusion $\nabla\cdot(k(x)\nabla u)$ or nonlinear advection $u\cdot\nabla u$, dimensional interactions become nontrivial due to cross-derivatives and nonlinear couplings. In such cases, decomposing the operator across active dimensions is less straightforward. While Anant-Net offers architectural flexibility to incorporate variable coefficients as inputs or represent nonlinear terms using network outputs and their gradients, doing so increases model complexity and requires careful redesign of the current architecture. These extensions remain the subject of ongoing work. Moreover, Anant-Net could also be extended to solve high-dimensional PDEs at selected points or along trajectories in the space rather than approximating the entire domain. To this end, we emphasize that solving PDEs in high dimensions over exponentially increasing domains is inherently difficult due to the curse of dimensionality, and Anant-Net aims to mitigate this challenge through a structured, scalable framework that lays the foundation for further generalizations.
Acknowledgement
We would like to thank all the reviewers for their valuable feedback and constructive suggestions, which have helped improve the quality and clarity of this work.
@@ Appendix A Universal Approximation Theorem: Through the lens of Anant-Net
(SPINN[29]) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$, we can find a sufficiently large $\tilde{r}>0$ and neural networks $f_{j}$ and $g_{j}$ such that,
$\biggl{\|}u-\sum_{j=1}^{\tilde{r}}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (15)
Theorem A.1 was the universal approximation theorem proposed by Cho et al.[29] in their work. However, as explained in the previous sections of this paper, naively implementing this approach for a high-dimensional partial differential equation is infeasible. To this end, we propose Anant-Net and the corresponding universal approximation theorem for Anant-Net, Theorem A.2.
Theorem A.2
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u^{k}\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$, we can find a sufficiently large $r>0$ and neural networks $f_{j}\triangleq f_{1,j}^{(\theta_{1})}$ and $g_{j}\triangleq f_{2,j}^{(\theta_{2})}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (16)
@@
where, $\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, $\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X\text{\textbackslash}\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y\text{\textbackslash}\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}.$
[…]
Proof. Let $\phi_{i}$} and $\psi_{j}$} be the orthonormal basis for $L^{2}(X)$ and $L^{2}(Y)$ respectively. Then $\phi_{i}\psi_{j}$} forms the orthonormal basis for $L^{2}(X\times Y)$. Therefore, we can find a sufficiently large r, consider an active dimension set $\mathcal{B}_{k}$ that results in the solution $u^{k}$ such that,
@@
$\|II_{1}\|^{2}_{L^{2}(X\times Y)}=\int_{X\times Y}\biggl{\{}\sum_{i,j}^{r}a_{i,j}^{k}\phi_{i}(\psi_{j}-f_{j})\biggr{\}}^{2}dxdy$
$\hskip 59.75095pt=\int_{X\times Y}\biggl{\{}\sum_{j}^{r}\biggl{(}\sum_{i}^{r}a_{i,j}^{k}\phi_{i}\biggr{)}(\psi_{j}-f_{j})\biggr{\}}^{2}dxdy$
$\hskip 17.07182pt=\int_{X\times Y}\biggl{\{}\sum_{j}^{r}A_{j}(x)B_{j}(y)\biggr{\}}^{2}dxdy$
@@
$\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\sum_{k=1}^{|\hat{\mathcal{B}}|}\varepsilon=|\hat{\mathcal{B}}|\cdot\varepsilon$
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}b_{i}\tilde{f_{i}}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}\{b_{i}\tilde{f_{i}}\}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}=\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}t_{i}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{r}f_{i}g_{i}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
@@ Appendix B Hyperparameters for Anant-Net
The main hyperparameters for Anant-Net include: the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#CG\#CG), number of boundary/data grids (#BG\#BG), and the sampling frequency ($S_{\text{FREQ}}$). The hyperparameters used for the main results presented in this paper are discussed in this section. For clarity, the tables presented below follow the format: #\#Collocation $(N_{C}\times N_{C}\times N_{C})\times\#CG$ and #\#Data $(N_{B}\times N_{B}\times N_{B})\times\#BG$ for the current cases dealing with 3 body networks. However, if any future work deals with 4 body networks, then #\#Collocation $(N_{C}\times N_{C}\times N_{C}\times N_{C})\times\#CG$ and #\#Data $(N_{B}\times N_{B}\times N_{B}\times N_{B})\times\#BG$ and likewise.
@@
Dimension Architecture Optimizer Sampling Frequency #\#Iterations
@@
Table 9: Anant-Net architecture used for solving high-dimensional Poisson equation.
Dimension #\#Collocation #\#Data
@@
Table 10: Details on collocation and data points sampled for cases in Table 9 for solving high-dimensional Poisson equation.
@@
Dimension Architecture Optimizer Sampling Frequency #\#Iterations
@@
Table 11: Anant-Net architecture used for solving high-dimensional sine-Gordon equation.
Dimension #\#Collocation #\#Data
@@
Table 12: Details on collocation and data points sampled for cases in Table 11 for solving high-dimensional sine-Gordon equation.
@@
Dimension Architecture Optimizer Sampling Frequency #\#Iterations
@@
Table 13: Anant-Net architecture used for solving high-dimensional Allen-Cahn equation.
Dimension #\#Collocation #\#Data
@@
Table 14: Details on collocation and data points sampled for cases in Table 13 for solving high-dimensional Allen-Cahn equation.
@@
Equation Dimension Architecture Optimizer Sampling Frequency #\#Iterations
@@
Table 15: Anant-KAN architecture used for solving high-dimensional Poisson equation, sine-Gordon equation and Allen-Cahn equation. Note that Spline basis is used for all the cases with polynomial order ($k$) = 2 and grid size ($G$) = 5.
Equation Dimension #\#Collocation #\#Data
@@
Hyperparameters for Section 3.6
Dimension Architecture Optimizer Sampling Frequency #\#Iterations
20D $\bigl{[}1,64,64,15\bigr{]}\times 1,\bigl{[}10,64,64,15\bigr{]}\times 2$ Adam 5000 100000
100D $\bigl{[}1,64,64,50\bigr{]}\times 1,\bigl{[}50,64,64,50\bigr{]}\times 2$ Adam 5000 300000
300D $\bigl{[}1,64,64,160\bigr{]}\times 1,\bigl{[}150,64,64,160\bigr{]}\times 2$ Adam 2500 375000
Table 17: Anant-Net architecture used for solving high-dimensional Heat equation.
Dimension #\#Collocation #\#Data
20D $(14\times 14\times 14)\times 20=54880$ $(10\times 10\times 10)\times 50=50000$
100D $(14\times 14\times 14)\times 60=164640$ $(10\times 10\times 10)\times 90=90000$
300D $(14\times 14\times 14)\times 150=411600$ $(10\times 10\times 10)\times 160=160000$
Table 18: Details on collocation and data points sampled for cases in Table 17 for solving high-dimensional Heat equation.
Dimension Architecture Optimizer Sampling Frequency #\#Iterations
100D $\bigl{[}1,5,5,18\bigr{]}\times 1,\bigl{[}50,5,5,18\bigr{]}\times 2$ Adam 1000 55000
Table 19: Anant-KAN architecture used for solving high-dimensional Heat equation. Note that Spline basis is used for all the cases with polynomial order ($k$) = 2 and grid size ($G$) = 5.
Dimension #\#Collocation #\#Data
100D $(14\times 14\times 14)\times 50=137200$ $(10\times 10\times 10)\times 60=60000$
Table 20: Details on collocation and data points sampled for cases in Table 19 for solving high-dimensional Heat equation.